OpenStudy (anonymous):

Which number can replace the variable to make these fractions equivalent n/9 = 56/63

OpenStudy (anonymous):

Cross multiply n * 63 9 * 56

OpenStudy (anonymous):

9*56=504

OpenStudy (anonymous):

n*63=63n???

OpenStudy (anonymous):

63/504=8

OpenStudy (anonymous):

Yep, you got it 8 is correct.